OpenStudy (anonymous):

simplify 3 square root 5 x 3 square root 40. Show work & don't leave answer in decimal

OpenStudy (rulnick):

3 sqrt(5) 3 sqrt(40) = 9 sqrt(5*5*2*2*2) = 9*5*2sqrt(2) = 90 sqrt(2)